OpenStudy (dls):

lim x->infinity (sinx+cosx)/(x^2)=?

OpenStudy (dls):

\[\LARGE \frac{(1-\frac{x^2}{2}+\frac{x^4}{24})+(x-\frac{x^3}{6}+\frac{x^5}{120})}{x^2}\] Tried using series..I got 1/2 i guess.

OpenStudy (goformit100):

it's Easy, i think this question is repeated by you.

OpenStudy (anonymous):

try to remove x from denomiator

OpenStudy (anonymous):

then, replace x with 0

OpenStudy (dls):

i get -1/2

OpenStudy (goformit100):

Good

OpenStudy (dls):

wrong

OpenStudy (dls):

Summons @ParthKohli

OpenStudy (dls):

Summon successful.

OpenStudy (dls):

Summons @terenzreignz

OpenStudy (dls):

Summon successful.

terenzreignz (terenzreignz):

Just for that, I'm fanning you. Now, stand by... while I assess the situation.

OpenStudy (dls):

I await sire.

terenzreignz (terenzreignz):

Okay, I have a verdict, but I think I'll wait for ParthKohli, he seems to be putting quite a bit of effort into this :D

OpenStudy (dls):

Is it so? @ParthKohli :O

OpenStudy (dls):

Why is -1/2 wrong though ?

terenzreignz (terenzreignz):

You'll see. The limit is zero, by the way :)

OpenStudy (dls):

hmm..

Parth (parthkohli):

Hey, guys I don't know how to do this one :-)

OpenStudy (dls):

LOL.

Parth (parthkohli):

JK hahahaha! :-P

terenzreignz (terenzreignz):

Haha, @ParthKohli L'Hopital doesn't apply, the numerator doesn't go to infinity :P

OpenStudy (dls):

Exactly.

Parth (parthkohli):

*close enough*

terenzreignz (terenzreignz):

My turn?

OpenStudy (dls):

Nevermind.

OpenStudy (dls):

But why is series going wrong?!

Parth (parthkohli):

I'll just keep typing and testing, don't expect much from me in the next few minutes.

terenzreignz (terenzreignz):

I have no idea, but it takes a clever bit of analysis, to realise... \[\huge -\frac{2}{x^2}<\frac{\sin x+\cos x}{x^2}<\frac2{x^2}\]And SQUEEZE it :P Both -2/x^2 and 2/x^2 go to zero :P

OpenStudy (dls):

\[\LARGE \cancel x^2\frac{(\frac{1}{x^2}-\frac{1}{2}+\frac{x^2}{24})+(\frac{1}{x^2}-\frac{x}{6}+\frac{x^3}{120})}{\cancel x^2}\] other terms with 1/x goes to 0..we are left with -1/2!!

Parth (parthkohli):

\[\dfrac{\sin(x)}{x^2} + \dfrac{\cos(x)}{x^2} = \dfrac{1}{x} \left(\dfrac{\sin(x)}{x} + \dfrac{\cos(x)}{x}\right)\]Yoohoo, yoohoo.

OpenStudy (dls):

^^

terenzreignz (terenzreignz):

I just used Squeeze. Who needs L'hopital :P

OpenStudy (dls):

But why is series wrong..somenoe?!

terenzreignz (terenzreignz):

Better consult an expert with series... I'm no such expert :D

OpenStudy (dls):

hmm..okay I liked @ParthKohli 's solution :P seems more legit :D

terenzreignz (terenzreignz):

No matter :D As long as you agree that it's zero.

OpenStudy (dls):

\[\LARGE \frac{xtan2x-2xtanx}{(1-\cos2x)^2}\] x->0 anyone? :P

terenzreignz (terenzreignz):

Okay, L'Hopital looks good here.

OpenStudy (dls):

I have a doubt.

OpenStudy (dls):

We won't use quotient rule here right?

Parth (parthkohli):

The hell with the product rule. >_>

OpenStudy (dls):

Just differentiate them seperately?

terenzreignz (terenzreignz):

Nope, just differentiate the numerator and the denominator.

Parth (parthkohli):

@DLS You cancelled the \(x^2\)s? Where did the \(x^2\) in the numerator come from? o_O

terenzreignz (terenzreignz):

Factored out an x^2 from the series.

OpenStudy (dls):

^^

Parth (parthkohli):

Ah, thanks

OpenStudy (dls):

so i get 1/x terms which tend to 0.

OpenStudy (dls):

OH GUYS SORRY!! SERIES CANT BE USED HERE >.<

OpenStudy (dls):

Series are only applicable for x<1 Foolish me.!!

terenzreignz (terenzreignz):

Only two of them, you still get terms with x^n in the numerators, some positive, some negative. @DLS Why complicate limits with series? :P

Parth (parthkohli):

Stay within your LIMITS, or you'd get a SERIES of disasters!

OpenStudy (dls):

haha

OpenStudy (dls):

yaayyy! free maths help!!

terenzreignz (terenzreignz):

What's OS for, then? :/

Parth (parthkohli):

OS is for free math cheating!

terenzreignz (terenzreignz):

Way to promote the site, ambassasdor ^ :D

Parth (parthkohli):

^_^

OpenStudy (dls):

\[\LARGE \frac{\tan2x+x^2\sec2x-2tanx-2xsec^2x}{2(1-\cos2x) \times?}\] not sure how to do the denominator :/ not sure about numerator either :P

Parth (parthkohli):

http://wolframalpha.com

OpenStudy (anonymous):

is this a limits again?

OpenStudy (dls):

16sin^3xcosx how :/

terenzreignz (terenzreignz):

\[\huge \frac{d}{dx}(1-\cos 2x)=2\sin2x\]

Parth (parthkohli):

^ Chain Rule :')

OpenStudy (yrelhan4):

The answer would be 1/2 when x-->0

OpenStudy (dls):

but its not so

Parth (parthkohli):

:-O RnR?

OpenStudy (yrelhan4):

I was bored man. @ParthKohli

terenzreignz (terenzreignz):

I'm going to go for a bit again... be back later, maybe ^.^ --------------------------------------------------- Terence out

OpenStudy (dls):

:(

Parth (parthkohli):

):

OpenStudy (dls):

How do i get the damn answer?

Parth (parthkohli):

Chain Rule aata hai?

OpenStudy (dls):

haan

OpenStudy (dls):

\[\LARGE \frac{x^2secx-2xsec^2x}{2(1-\cos2x) \times (2\sin2x) \times (1-\tan^2x)}\] im left with this pellet :P

Parth (parthkohli):

:-O

OpenStudy (dls):

??

OpenStudy (dls):

i changed tan2x to its formula,so i can eliminate it

OpenStudy (yrelhan4):

Differentiate mat karo.. tan2x ko kholo=(2tanx/1-tan^2(x)) 2xtanx ko bahar nikal lo. Bach jayega [2xtanx(tan^2(x)]/[4sin^4x(1-tan^2(x)] Isko cos aur sin me convert karo.. 1/2[xtanx]/sin^2x*(cos^2x-sin^2x). Isko solve kro 1/2 aayega.

OpenStudy (dls):

galaat hai to kyun padu be :/

OpenStudy (yrelhan4):

Bhaad me jaa.

OpenStudy (dls):

:)